Reports 1-1 of 1 Clear search Modify search
AdV-TCS (CO2 laser projector)
taranto, bersanetti, pinto - 23:18 Friday 09 February 2024 (63222) Print this report
CO2 DAS power tuning @18 W, pt.5

The afternoon shift started from the CO2 powers configuration left in the previous shift (entry 63219) and reported in the following table:


CH [W]




on the ITF




on the pickoff 





on the ITF




on the pickoff 




We tested the lock acquisition of this CO2 powers configuration and at 14:50 UTC LN2 was reached. After 40 minutes (at 15:31 UTC), the WI DAS OUT was increased by +50 mW (1.52 W 1.57 W), but at 16:42 UTC, the ITF unlocked (see dedicated entry). In Figure 1, the main ITF signals of this lock are shown, no big impact has been observed.

From this moment, it was easy to lock the CITF, but the value of the sidebands was low, the 12 MHz did not even exceed 0.05 mW. To address this situation, the NI CH was increased in order to follow the increase of the NI DAS OUT made in the morning. Two steps were taken: the power of the NI CH was increased by 1 mW (0.055 W 0.056 W) at 18:29 UTC and then by 2 mW (0.056 W 0.058 W) at 20:27 UTC, but again, no good effect was observed, see Figure 2.

Given the unclear situation, we decided to return to the starting point of the afternoon (where the sidebands at CITF were good and we had reached LN2). Therefore, at 20.51 UTC we reverted the NI CH and the WI DAS OUT back to their values of the starting point (NI CH: 0.055 W and WI DAS OUT: 1.52 W). In this case, however, the situation doesn't seem to have changed much either.

Images attached to this report
Comments to this report:
bersanetti, mpinto, taranto, spinicelli - 23:29 Friday 09 February 2024 (63224) Print this report

From the ISC side, along the shift we didn't tune any of our locking parameters, but monitored the different TCS steps.

Here we report the most relevant notes of the shift:

  • The unlock of 16.40 UTC was due to a 3.5 mag earthquake nearby (Parma). This again caused the EIB bench to move a lot, with the BPC trying to catch up with the movement of the bench. The PMC safety triggered putting the PMC in scan mode with consequent unlock of the injection system. Together with Piernicola, the situation was recovered and we kept with the thermal tuning.
  • Also today we experienced the same issue of the past days concerning the B2 QD2 GALVO correction being too high, see Fig.1. We noticed, besides the actions performed the thermal tuning, that the corrections were varying a lot also during static conditions. (We are not sure about the meaning of this behaviour, but we report anyhow the plot with the occurrence).
  • At around 22.10 UTC some series of bursts appeared on the 112 MHz and 12 MHz sidebands and on B1p, at the same time the EOM corrections started to increase. Few minutes later, we unlocked from the DRMI 3f state. (maybe related to the ALS).

We left the ITF locked on the IR.

Images attached to this comment
Search Help